### v2

@@ 4 Rectified exclusion core
The possibility for the strong core and the exclusion core to lack predictive power under weak preferences motivates us to analyze the underlying causes and seek potential rectifications. This leads to our proposal of the rectified exclusion core.
@@
We begin our analysis by examining the behavioral foundation for unaffected agents in a weak blocking coalition. The cooperative approach does not explain why an unaffected agent may be willing to join such a coalition. In the absence of side payments, a seemingly intuitive explanation is altruism: unaffected agents are willing to help the others in the coalition without incurring any cost to themselves. However, this so-called altruism is inherently biased: it benefits agents within the coalition but may harm those outside it. In Example 1, under $^{\prime}_{I}$, agent $2$ is willing to help either agent $1$ or agent $3$ and alternates between them in forming blocking coalitions, which results in an empty strong core.
@@
balbuzanov2019endowments’s concept of exclusion rights offers a more plausible explanation for the behavior of unaffected agents. Consider strict preferences. Suppose that a coalition $C$ weakly blocks an allocation $\mu$ via another $\sigma$, and $C_{\sigma\sim\mu}\neq\emptyset$. Therefore, for every $i\in C_{\sigma\sim\mu}$, $\sigma(i)=\mu(i)$. If there exists a subcoalition $C^{\prime}\subseteq C_{\sigma\sim\mu}$ such that $\mu(C^{\prime})=\omega(C^{\prime})$, let $C^{\prime}$ be the largest subcoalition satisfying this condition. [^6] This largest $C^{\prime}$ must be unique, because for any two subcoalitions $C^{a}$ and $C^{b}$ such that $\mu(C^{a})=\omega(C^{a})$ and $\mu(C^{b})=\omega(C^{b})$, we have $\mu(C^{a}\cup C^{b})=\omega(C^{a}\cup C^{b})$. So, $C^{\prime}$ does not contribute objects to the others in the coalition under both $\mu$ and $\sigma$. Thus, they can removed from the coalition without any loss. For every $i\in C_{\sigma\sim\mu}\backslash C^{\prime}$, we must be able to find a chain, denoted by $i\rightarrow i_{1}\rightarrow i_{2}\rightarrow\cdots\rightarrow j$, in which every agent receives the next agent’s endowment under $\mu$ and the end of the chain is an agent from $C_{\sigma>\mu}$. This means that the objects received by $\CC_{\sigma\sim\mu}\backslash C^{\prime}$ in $\mu$ are directly or indirectly controlled by $C_{\sigma>\mu}$. Therefore, the participation of $\CC_{\sigma\sim\mu}\backslash C^{\prime}$ in the coalition can be viewed as the result of threat exerted by $C_{\sigma>\mu}$, who hold exclusion rights over $\CC_{\sigma\sim\mu}\backslash C^{\prime}$. In fact, this is the argument used by balbuzanov2019endowments to prove the equivalence between the strong core and the exclusion core in the housing market model under strict preferences.
Since exclusion rights can effectively explain coalition formation in standard core concepts under strict preferences, we are motivated to examine exclusion rights and the associated concept of the exclusion core under weak preferences.
@@
The idea behind the exclusion core is that agents have the right to evict others from the objects they control. Our first observation is that, under weak preferences, the definition of exclusion rights needs to be reformulated. Regardless of whether preferences are strict or not, it is natural that every coalition $C$ directly controls their endowments $\omega(C)$ and has the right to evict any others from $\omega(C)$ in any allocation. Under strict preferences, $C$ can indirectly control the endowments of the agents who occupy $\omega(C)$ by threatening to evict them from $\omega(C)$. This threat is credible because those agents cannot find other objects as good as $\omega(C)$. Under weak preferences, however, this threat might become incredible, because those agents may find other objects as good as $\omega(C)$. For example, if an agent is indifferent between all objects, then he is not fear of any threat because he can remain unaffected by consuming his own endowment. Therefore, in our new definition, $C$ indirectly controls an agent $i$’s endowment in an allocation $\mu$ if, not only $\mu(i)$, but also all the objects that $i$ views as good as $\mu(i)$ are controlled by $C$.
@@
where $C^{0}=C$, and $C^{k}=C^{k-1}\cup\{i\in I\backslash C^{k-1}:\mathcal{I}_{i}(\mu(i))\subseteq\omega(C^{k-1})\}$ for every $k\geq 1$. Under strict preferences, $^{*}(C|\omega,\mu)$ reduces to the definition used by balbuzanov2019endowments.
@@
The following two examples illustrate the scenarios in which we allow unaffected agents to join exclusion blocking coalitions.
@@
Consider the following market with three agents. For convenience, we use a graph to indicate each agent’s favorite object. This convention is followed by other examples in this section.
@@
Since agent $1$ most prefers agent $2$’s endowment, and $2$ views $1$’s endowment as one of his favorite objects but does not accept agent $3$’s endowment, the intuitive allocation in this market is $\sigma$, in which $1$ and $2$ exchange endowments. The strong core = the exclusion core = $\{\sigma\}$.
@@
Nevertheless, if we examine the market more carefully, we will see that $1$ can actually achieve the goal of evicting $3$ from $b$ by exercising his direct exclusion right. Specifically, since $1$ directly controls $a$, if $1$ evicts $2$ from $a$, 22 must reclaim his endowment $b$ to remain unaffected. Then, $3$ must be evicted from $b$.
Example 4 illustrates one scenario in which we allow unaffected agents to join an exclusion blocking coalition: $\bigcupop\displaylimits_{i\in C_{\sigma\sim\mu}}\mathcal{I}_{i}(\mu(i))\backslash\omega(C_{\sigma\sim\mu})\subseteq\omega(C_{\sigma>\mu})$, and there exists a one-to-one mapping $g:C_{\sigma\sim\mu}\rightarrow\omega(C_{\sigma\sim\mu})$ such that for every $i\in C_{\sigma\sim\mu}$, $g(i)\sim_{i}\mu(i)$.
In other words, $C_{\sigma>\mu}$ does not control the endowments of $C_{\sigma\sim\mu}$, because $C_{\sigma\sim\mu}$ can remain unaffected by allocating their endowments among themselves. However, aside from their endowments, all other objects that make $C_{\sigma\sim\mu}$ unaffected are owned by $C_{\sigma>\mu}$. Thus, if $C_{\sigma>\mu}$ evicts $C_{\sigma\sim\mu}$ from $\omega(C_{\sigma>\mu})$, $C_{\sigma\sim\mu}$ must revert to $\omega(C_{\sigma\sim\mu})$ to remain unaffected, which must evict others from $\omega(C_{\sigma\sim\mu})$. Thus, $C_{\sigma>\mu}$ can indirectly evict others from $\omega(C_{\sigma\sim\mu})$ by exercising their direct exclusion rights. To demonstrate this exclusion right, we allow $C_{\sigma\sim\mu}$ and $C_{\sigma>\mu}$ to jointly form a coalition. In Example 4, we allow $1$ and $2$ to form a coalition to evict $3$ from $b$.
@@
  • If $3$ can successfully obtain $b$ after being evicted from $d$, 11 will be evicted from $b$, which is his unique favorite object. Then, $4$ can leverage this indirect threat to obtain the control right of $1$’s endowment $a$, and then evict $5$.
@@
  • If $3$ cannot obtain $b$ after being evicted from $d$, 44’s threat towards $3$ becomes credible. Leveraging this threat, $4$ can obtain the control right of $c$ and threaten to evict $2$ from $c$. To counteract this threat, $2$ will need to obtain $a$ to remain unaffected. Similarly as above, if $2$ can successfully obtain $a$, 55 will be evicted from $a$. If $2$ cannot obtain $a$, 44’s indirect threat towards $2$ becomes credible, and then $4$ can obtain the control right of $b$. Then, $4$ can threaten $1$ to obtain the control right of $a$, and then evict $5$.
@@
Example 5 illustrates another scenario in which we allow unaffected agents to join an exclusion blocking coalition: although $C_{\sigma>\mu}$ does not control the endowments of $C_{\sigma\sim\mu}$, we have $\bigcupop\displaylimits_{i\in C_{\sigma\sim\mu}}\mathcal{I}_{i}(\mu(i))\backslash\omega(C_{\sigma\sim\mu})\subseteq\omega(C_{\sigma>\mu})$, and there does not exist a one-to-one mapping $g:C_{\sigma\sim\mu}\rightarrow\omega(C_{\sigma\sim\mu})$ such that, for every $i\in C_{\sigma\sim\mu}$, $g(i)\sim_{i}\mu(i)$.
@@
In other words, $C_{\sigma\sim\mu}$ must rely on the endowments of $C_{\sigma>\mu}$ to remain unaffected. By leveraging the threat of evicting $C_{\sigma\sim\mu}$ from $\omega(C_{\sigma>\mu})$, $C_{\sigma>\mu}$ can indirectly control $\omega(C_{\sigma\sim\mu})$ and pressure $C_{\sigma\sim\mu}$ to join the coalition. In Example 5, $\{1,2,3\}$ relies on $4$’s endowment to remain unaffected. Thus, we allow $\{1,2,3,4\}$ to form a coalition to exclusion block $\mu$.
@@
To unify the above scenarios, our definition imposes the following requirement on an exclusion blocking coalition $C$: if $C_{\sigma\sim\mu}\neq\emptyset$, then $\forall i\in C_{\sigma\sim\mu}$, $\mathcal{I}_{i}(\mu(i))\backslash\omega(C_{\sigma\sim\mu})\subseteq\omega(C_{\sigma>\mu})$, or equivalently, $\mathcal{I}_{i}(\mu(i))\subseteq\omega(C)$. In words, all the objects that make every $i\in C_{\sigma\sim\mu}$ unaffected are owned by $C$. In this case, if $C_{\sigma>\mu}$ evicts $C_{\sigma\sim\mu}$ from $\omega(C_{\sigma>\mu})$, then, either $C_{\sigma\sim\mu}$ can revert to their own endowments to remain unaffected, which evicts others from $\omega(C_{\sigma\sim\mu})$, or $C_{\sigma\sim\mu}$ cannot rely on their own endowments to remain unaffected, thereby justifying their participation in the coalition. [^7] Our definition accommodates the case that, for every $i\in C_{\sigma\sim\mu}$, $\mathcal{I}_{i}(\mu(i))\subseteq\omega(C_{\sigma>\mu})$. So, $C_{\sigma>\mu}$ indirectly controls the endowments of $C_{\sigma\sim\mu}$. In this case, we have flexibility to define the blocking coalition: either $C_{\sigma>\mu}$ alone forms a coalition to evict others from $\omega(C_{\sigma\sim\mu})$, or $C_{\sigma>\mu}$ and $C_{\sigma\sim\mu}$ jointly form a coalition to evict others from $\omega(C_{\sigma\sim\mu})$. A difference is that, if $C_{\sigma\sim\mu}$ does not join the blocking coalition, they are allowed to be harmed by the blocking, but if they join the blocking coalition, they cannot be harmed.
@@
The rectified exclusion core consists of allocations that are not rectifyingly exclusion blocked.
Our main result in this section is the following theorem.
@@
(1) The rectified exclusion core is nonempty, Pareto efficient, and equivalence-closed. (2) It is a superset of the exclusion core and a subset of the weak core. (3) It coincides with the strong core (and with the exclusion core) when the strong core core is nonempty.
We remark on the equivalence-closedness property. As shown in Proposition 2, the strong core and the weak core satisfy this property, whereas the exclusion core violates it. The issue with the exclusion core is that a coalition $C$ may have different exclusion rights in equivalent allocations. After reformulating exclusion rights, we ensure that $^{*}(C|\omega,\mu)={}^{*}(C|\omega,\mu^{\prime})$ whenever $\mu$ and $\mu^{\prime}$ are equivalent. As a result, the rectified exclusion core satisfies the property.

### v3

@@ 4 Rectified exclusion core
The limited predictive power of the strong core and of the exclusion core under weak preferences motivates us to analyze the underlying causes and seek rectifications.
@@
We begin our analysis by examining the behavioral foundation of coalition formation under the strong core. A weak blocking coalition may include agents who are unaffected by the blocking. The cooperative approach does not explain why unaffected agents are willing to join blocking coalitions. Without transfers, a seemingly intuitive explanation is that agents are willing to help others without incurring any cost to themselves. However, this “altruism” interpretation is flawed, since unaffected agents help those within a blocking coalition but may harm outside agents. In Example 1, under $^{\prime}_{I}$, agent $2$ may help either agent $1$ or agent $3$ and alternate between them in forming blocking coalitions, rending the strong core empty.
@@
In contrast, BK’s concept of exclusion rights offers a more plausible explanation for the behavior of unaffected agents. Consider strict preferences. Suppose that a coalition $C$ weakly blocks an allocation $\mu$ via another $\sigma$, and $C_{\sigma\sim\mu}\neq\emptyset$. So, $\sigma(C)=\omega(C)$, and for every $i\in C_{\sigma\sim\mu}$, $\sigma(i)=\mu(i)$. If there exists a subcoalition $C^{\prime}\subseteq C_{\sigma\sim\mu}$ such that $\mu(C^{\prime})=\omega(C^{\prime})$, let $C^{\prime}$ be the largest subcoalition satisfying this condition. [^7] This largest $C^{\prime}$ is unique, because for any two subcoalitions $C^{a}$ and $C^{b}$ such that $\mu(C^{a})=\omega(C^{a})$ and $\mu(C^{b})=\omega(C^{b})$, we have $\mu(C^{a}\cup C^{b})=\omega(C^{a}\cup C^{b})$. We view $C^{\prime}$ as redundant in the coalition, because $C^{\prime}$ does not contribute objects to the others in the coalition under both $\mu$ and $\sigma$. So, $C^{\prime}$ can be removed from $C$ without any loss. For every $i_{1}\in C_{\sigma\sim\mu}\backslash C^{\prime}$, we must be able to find a chain denoted by
$i_{1}\rightarrow i_{2}\rightarrow\cdots i_{K}\rightarrow j,$
in which $\{i_{1},\ldots,i_{K}\}\subseteq C_{\sigma\sim\mu}$, every $i_{k}$ receives $i_{k+1}$’s endowment under $\mu$, and the last agent $i_{K}$ receives the endowment of some $j\in C_{\sigma>\mu}$ under $\mu$. So, this means that the objects received by $\CC_{\sigma\sim\mu}\backslash C^{\prime}$ under $\mu$ are directly or indirectly controlled by $C_{\sigma>\mu}$. Then, the participation of $\CC_{\sigma\sim\mu}\backslash C^{\prime}$ in the blocking coalition can be viewed as the result of pressure due to $C_{\sigma>\mu}$, who hold exclusion rights over $\CC_{\sigma\sim\mu}\backslash C^{\prime}$. In fact, this is the argument used by BK to prove the coincidence of the strong core and the exclusion core in the housing market model under strict preferences.
In the following, we examine how to develop the idea of exclusion rights to define coalition formation under weak preferences. This leads to our modification of the exclusion core in this section and the modification of the strong core in the next section.
@@
The idea behind the exclusion core is that agents have the right to evict others from the objects they control. Our first observation is that, under weak preferences, the definition of indirect control needs to be reformulated. Regardless of strict or weak preferences, in any allocation, it is natural that every coalition $C$ directly controls their endowments $\omega(C)$ and has the right to evict any others from $\omega(C)$. Under strict preferences, BK argue that $C$ can indirectly control the endowments of the agents who occupy $\omega(C)$ by threatening to evict them from $\omega(C)$. This threat is credible because those agents cannot find other objects as good as $\omega(C)$. Under weak preferences, however, this threat might become incredible, because those agents may find other objects as good as $\omega(C)$. For example, if an agent is indifferent between all objects, then he is not fear of any threat because he can remain unaffected by consuming his own endowment. So, in our new definition, $C$ indirectly controls an agent $i$’s endowment in an allocation $\mu$ if not only $\mu(i)$ but also all the objects that $i$ views as good as $\mu(i)$ are controlled by $C$.
@@
where $C^{0}=C$, and $C^{k}=C^{k-1}\cup\{i\in I\backslash C^{k-1}:\mathcal{I}_{i}(\mu(i))\subseteq\omega(C^{k-1})\}$ for every $k\geq 1$. Under strict preferences, $^{*}(C|\omega,\mu)$ reduces to the definition used by BK.
@@
Nevertheless, we observe that under weak preferences, the execution of exclusion rights also becomes subtle, requiring unaffected agents to join exclusion blocking coalitions in certain scenarios. The following two examples illustrate those scenarios.
@@
Consider the following market with three agents. For convenience, we use a graph to indicate each agent’s favorite object. This convention is followed by remaining examples.
@@
Since $1$ most prefers $2$’s endowment and $2$ views $1$’s endowment as one of his favorite objects and prefers his endowment over $3$’s, the allocation $\sigma$, in which $1$ and $2$ exchange endowments, is the only intuitive allocation in the market. The strong core = the exclusion core = $\{\sigma\}$.
@@
Nevertheless, if we examine the market more carefully, we will see that $1$ can achieve the goal of evicting $3$ from $b$ by exercising his direct exclusion right. Specifically, since $1$ directly controls $a$, if $1$ evicts $2$ from $a$, 22 must reclaim his endowment $b$ to maintain his welfare. Then, $3$ must be evicted from $b$.
Example 4 illustrates one scenario in which we allow unaffected agents to join an exclusion blocking coalition: $\bigcupop\displaylimits_{i\in C_{\sigma\sim\mu}}\mathcal{I}_{i}(\mu(i))]\backslash\omega(C_{\sigma\sim\mu})\subseteq\omega(C_{\sigma>\mu})$ and there exists a one-to-one mapping $g:C_{\sigma\sim\mu}\rightarrow\omega(C_{\sigma\sim\mu})$ such that for every $i\in C_{\sigma\sim\mu}$, $g(i)\sim_{i}\mu(i)$.
In other words, $C_{\sigma>\mu}$ does not control the endowments of $C_{\sigma\sim\mu}$, because $C_{\sigma\sim\mu}$ can remain unaffected by reallocating their endowments among themselves. However, aside from their endowments, all other objects that make $C_{\sigma\sim\mu}$ unaffected are owned by $C_{\sigma>\mu}$. So, if $C_{\sigma>\mu}$ evicts $C_{\sigma\sim\mu}$ from $\omega(C_{\sigma>\mu})$, to remain unaffected, $C_{\sigma\sim\mu}$ must revert to $\omega(C_{\sigma\sim\mu})$ and evict others from $\omega(C_{\sigma\sim\mu})$. So, $C_{\sigma>\mu}$ can indirectly evict others from $\omega(C_{\sigma\sim\mu})$ by exercising their direct exclusion rights. To demonstrate these exclusion rights, we allow $C_{\sigma\sim\mu}$ and $C_{\sigma>\mu}$ to jointly form a blocking coalition. In Example 4, $1$ and $2$ can form a coalition to evict $3$ from $b$.
@@
  • If $3$ obtains $b$ after being evicted from $d$, 11 will be evicted from $b$, which is his unique favorite object. Then, $4$ can leverage this indirect threat to obtain the control right of $1$’s endowment, and then evict $5$.
@@
  • If $3$ cannot obtain $b$ after being evicted from $d$, 44’s threat towards $3$ becomes credible. Leveraging this threat, $4$ can obtain the control right of $c$ and threaten to evict $2$ from $c$. To counteract this threat, $2$ will need to obtain $a$ to remain unaffected. Similarly as above, if $2$ obtains $a$, 55 will be evicted from $a$. If $2$ cannot obtain $a$, 44’s indirect threat towards $2$ becomes credible, and then $4$ can obtain the control right of $b$. Then, $4$ can threaten $1$ to obtain the control right of $a$, and then evict $5$.
@@
Example 5 illustrates another scenario in which we allow unaffected agents to join an exclusion blocking coalition: while $\bigcupop\displaylimits_{i\in C_{\sigma\sim\mu}}\mathcal{I}_{i}(\mu(i))\nsubseteq\omega(C_{\sigma>\mu})$, it holds that $\bigcupop\displaylimits_{i\in C_{\sigma\sim\mu}}\mathcal{I}_{i}(\mu(i))]\backslash\omega(C_{\sigma\sim\mu})\subseteq\omega(C_{\sigma>\mu})$ and there does not exist a one-to-one mapping $g:C_{\sigma\sim\mu}\rightarrow\omega(C_{\sigma\sim\mu})$ such that, for every $i\in C_{\sigma\sim\mu}$, $g(i)\sim_{i}\mu(i)$.
@@
In other words, although $C_{\sigma>\mu}$ does not control the endowments of $C_{\sigma\sim\mu}$, $C_{\sigma\sim\mu}$ must rely on the endowments of $C_{\sigma>\mu}$ to remain unaffected. In this case, we allow $C_{\sigma>\mu}$ to pressure $C_{\sigma\sim\mu}$ to join the coalition by threatening to evict $C_{\sigma\sim\mu}$ from $\omega(C_{\sigma>\mu})$. This justifies the participation of $C_{\sigma\sim\mu}$ in $C$. In Example 5, $\{1,2,3\}$ relies on $4$’s endowment to remain unaffected. Thus, we allow they form a coalition to exclusion block $\mu$.
@@
To unify the above scenarios, our definition imposes the following requirement on an exclusion blocking coalition $C$: if $C_{\sigma\sim\mu}\neq\emptyset$, then for all $i\in C_{\sigma\sim\mu}$, $\mathcal{I}_{i}(\mu(i))\backslash\omega(C_{\sigma\sim\mu})\subseteq\omega(C_{\sigma>\mu})$. This condition is equivalent to requiring that, for all $i\in C_{\sigma\sim\mu}$, $\mathcal{I}_{i}(\mu(i))\subseteq\omega(C)$. In words, all the objects that make every $i\in C_{\sigma\sim\mu}$ unaffected are owned by $C$. So, if $C_{\sigma>\mu}$ evicts $C_{\sigma\sim\mu}$ from $\omega(C_{\sigma>\mu})$, then, either $C_{\sigma\sim\mu}$ can revert to their own endowments to remain unaffected, which must evict others from $\omega(C_{\sigma\sim\mu})$, or $C_{\sigma\sim\mu}$ must rely on the endowments of $C_{\sigma>\mu}$ to remain unaffected, justifying their participation in the coalition. [^8] Our definition accommodates the case that, for every $i\in C_{\sigma\sim\mu}$, $\mathcal{I}_{i}(\mu(i))\subseteq\omega(C_{\sigma>\mu})$. So, $C_{\sigma>\mu}$ indirectly controls the endowments of $C_{\sigma\sim\mu}$. In this case, we have flexibility to define the blocking coalition: either $C_{\sigma>\mu}$ alone forms a coalition to evict others from $\omega(C_{\sigma\sim\mu})$, or $C_{\sigma>\mu}$ and $C_{\sigma\sim\mu}$ jointly form a coalition to evict others from $\omega(C_{\sigma\sim\mu})$. A difference between them is that, if $C_{\sigma\sim\mu}$ does not join the blocking coalition, they are allowed to be harmed by the blocking, but if they join the blocking coalition, they cannot be harmed.
@@
The rectified exclusion core consists of allocations that are not rectification exclusion blocked.
We apply this concept to previous examples.
Example revisited.
In $^{\prime}_{I}$ of Example 1, the rectified exclusion core = the exclusion core = $\{\mu,\delta\}$. Because $2$ is indifferent between all objects, the only possible blocking coalition he may join is the grand coalition $I$. Since $\mu$ or $\delta$ are Pareto efficient, they cannot be blocked.
In Example 2, the rectified exclusion core = $\{\mu,\sigma\}$, whereas the exclusion core = $\emptyset$. Agent $3$ cannot rectification exclusion block $\mu$ via $\sigma$, because $3$ does not indirectly control $2$’s endowment according to our definition, given $2$ is indifferent between $a$ and $c$. Similarly, $1$ cannot rectification exclusion block $\sigma$ via $\mu$.
In Example 3, the rectified exclusion core = $\{\mu,\sigma,\delta,\eta\}$, whereas the exclusion core = $\{\mu,\sigma\}$. Similar to Example 2, $3$ cannot rectification exclusion block $\delta$ via $\mu$ and $1$ cannot rectification exclusion block $\eta$ via $\sigma$, because they do not indirectly control $2$’s endowment.
In Example 4, the rectified exclusion core = the exclusion core = the strong core = $\{\sigma\}$. Allocation $\mu$ is rectification exclusion blocked by $\{1,2\}$ via $\sigma$.
In Example 5, the rectified exclusion core = the exclusion core = the strong core = $\{\sigma\}$. Allocation $\mu$ is rectification exclusion blocked by $\{1,2,3,4\}$ via $\sigma$.
Our main result in this paper is the following theorem.
@@
(1) The rectified exclusion core is nonempty, Pareto efficient, and equivalence-closed. (2) It is a superset of the exclusion core and a subset of the weak core. (3) It coincides with the strong core when the strong core core is nonempty.
As the exclusion core, every element of the rectified exclusion core is Pareto efficient, since every inefficient allocation can be blocked by the grand coalition $I$ via a Pareto improvement. After reformulating exclusion rights, we ensure that every coalition controls the same set of objects under equivalent allocations. As a result, the rectified exclusion core is equivalence-closed, which is in contrast to the exclusion core.
At first glance, it is unclear whether the rectified exclusion core is stronger or weaker than the original exclusion core, since we impose a more stringent condition on indirect exclusion but relaxes the formation of blocking coalitions by allowing unaffected agents to join. Theorem 1 proves that the rectified exclusion core is a superset of the exclusion core. While it is always nonempty, it is not overly large. It is a subset of the weak core and, importantly, coincides with the strong core when the strong core is nonempty.
Proposition 1 has proved that the exclusion core is a superset of the strong core. So, when the strong core is nonempty, the exclusion core also coincides with the strong core.
An algorithm provided in the next subsection can find elements of the rectified exclusion core, proving its nonemptiness.